Determine the value of the equilibrium constant (Kp), given the equilibrium partial pressures below:

NH4Br(s) ↔ NH3(g) + HBr(g)        

P of NH3 = 0.344 atm

P of HBr = 0.628 atm

Express your answer without units to 3 significant figures.
Answer #1

NH4Br(s) = NH3(g) + HBr(g)

Kp = P(NH3) * P(HBr)

Solid will not considered in equilibrium expression

= (0.344) * ( 0.628)

= 0.216

If you have any questions please comment

If you satisfied with the solution please rate it thanks